HardcoreTime100:

Solve this linear equation for p: 2.6(5.5p – 12.4) = 127.92 1. Distributive property: 2. Addition property of equality: 3. Division property of equality: 4. Solution: Solve this linear equation for p: 2.6(5.5p – 12.4) = 127.92 1. Distributive property: 2. Addition property of equality: 3. Division property of equality: 4. Solution: 14.3p – 32.24 = 127.92 14.3p – 32.24 + 32.24 = 127.92 + 32.24 14.3p = 160.16 StartFraction 14.3 p Over 14.3 EndFraction equals StartFraction 160.16 Over 14.3 EndFraction. p = 14.3p – 32.24 = 127.92 14.3p – 32.24 + 32.24 = 127.92 + 32.24 14.3p = 160.16 StartFraction 14.3 p Over 14.3 EndFraction equals StartFraction 160.16 Over 14.3 EndFraction. p =

Mercury:

first one: Solve this linear equation for p: 2.6(5.5p – 12.4) = 127.92 1. Distributive property: distribute the 2.6 across the parentheses 2.6*(5.5p) - 2.6(12.4) = 14.3 p - 32.24 = 127.92 2. Addition property of equality: add 32.24 to both sides to isolate 14.3p 3. Division property of equality: divide both sides by 14.3 to isolate p 4. Solution: report the result